OpenStudy (anonymous):

To what does the term i2 simplify?

ganeshie8 (ganeshie8):

\[\huge i = \sqrt{-1}\]

OpenStudy (anonymous):

thats the answer ?

OpenStudy (diyadiya):

\[\large i^2 = (\sqrt{-1})^2\]

OpenStudy (diyadiya):

since \[\large i = \sqrt{-1}\] as ganeshie8 said so what is \( \large i^2 \) ?